OpenStudy (anonymous):

A geometric series has a sum of 9,837. The common ratio is 3 and the first term is 9. How many terms are in the series?

OpenStudy (anonymous):

https://en.wikipedia.org/wiki/Geometric_series

OpenStudy (nincompoop):

this is fun

OpenStudy (nincompoop):

to find any term in geometric sequence \[a_n=a_1\times r^{n-1}\] a1 = first term r= common ratio n = number of the term to find

OpenStudy (nincompoop):

9837 = 9 * 3^(n-1) do you know logarithm?

OpenStudy (anonymous):

Yeah I would know how to find it out using the equation you wrote but I dont know the number of terms there are. that is what i am trying to find

OpenStudy (anonymous):

They give me multiple choice answers: A. 5 B. 7 C. 9 D. 11

OpenStudy (nincompoop):

if you know how to find out, then why can't you determine the terms?

OpenStudy (anonymous):

because the equation you wrote requires number of terms I am trying to find... lol sorry if I am being real dumb

OpenStudy (ranga):

\[\Large Sum = a \times \frac{ 1 - r^n }{ 1-r }\]

OpenStudy (nincompoop):

you can try plugging in and evaluate the sum

OpenStudy (ranga):

\[9837 = 9 \times \frac{ 1 - 3^n }{ 1-3 }\] Solve for n.

OpenStudy (nincompoop):

@ranga she is stuck in the same predicament as before

OpenStudy (nincompoop):

you can use the choices and plug them one by one until the sum equals

OpenStudy (nincompoop):

my guess would be 7 ;)

OpenStudy (anonymous):

I tried just pluggin them all in but for some reason none of them work...

OpenStudy (nincompoop):

try ranga's sum of series

OpenStudy (anonymous):

is it 7????

OpenStudy (ranga):

\[9837 = 9 \times \frac{ 1 - 3^n }{ 1-3 } = 9 \times \frac{ 3^n - 1 }{ 3-1 } = \frac{ 9 }{ 2 } \times (3^n-1)\]\[3^n -1 = 9837 \times \frac{ 2 }{ 9 } = 2186 \\ 3^n = 2187\]

OpenStudy (ranga):

yes. n = 7.

OpenStudy (nincompoop):

yaaaaa even using the first formula, you'll get 7 as an estimated value

OpenStudy (anonymous):

thank you both!!!!!!

OpenStudy (nincompoop):

something's amiss

OpenStudy (ranga):

Since there are just 7 terms, I can list the entire geometric series here: 9, 27, 81, 243, 729, 2187, 6561. 7 terms total. Add them up and the sum will be 9837.